# The Belle II Collaboration

### v1

@@ Abstract
A sample of 365 fb-1 of $e^{+}e^{-}\to\Upsilon(4S)\to B\bar{B}$ data collected by the Belle II experiment is used to measure the partial branching fractions of charmless semileptonic $B$ meson decays and determine the magnitude of the Cabibbo-Kobayashi-Maskawa matrix element $V_{ub}$. Events containing a signal electron or muon $\ell$ and a fully reconstructed hadronic $B$ decay that constrains the signal kinematics are selected, while the rest of the event defines the hadronic system $X_{u}$ associated with the signal. To discriminate the signal from the 50-times larger background originating from CKM-favored semileptonic $B$ decays, a template fit is performed in both signal and control regions after applying an optimized selection. The partial branching fraction measured for lepton energies greater than 1 GeV in the signal $B$ meson rest frame is $\Delta\mathcal{B}(B\rightarrow X_{u}\ell\nu)=(1.54\pm 0.08\,{\rm(stat.)}\pm 0.12\,{\rm(syst.)})\times 10^{-3}$. From this measurement, using the Gambino, Giordano, Ossola, Uraltsev theoretical framework, $|V_{ub}|=(4.01\pm 0.19^{+0.07}_{-0.08})\times 10^{-3}$ is determined, where the uncertainties are experimental and theoretical, respectively. This value is consistent with the world average obtained from previous inclusive measurements. Different theoretical predictions and partial branching fractions measured in other phase-space regions, defined by additional selections on the $X_{u}$ and leptonic system masses, are also used to determine $|V_{ub}|$.
@@ I Introduction
Precise measurements of the Cabibbo-Kobayashi-Maskawa (CKM) matrix elements [1, 2] are crucial to test the quark weak interactions described by the standard model of particle physics. For three fermion generations the CKM matrix is unitary, which imposes strict relations between its elements. The precision of experimental tests of several of these relations is limited by the current uncertainty in the determination of the strength of the $b$-to-$u$ quark coupling, $|V_{ub}|$. $B$-meson decays to either an electron or a muon, $\ell$, a neutrino, $\nu$, and a charmless hadronic system, $X_{u}$, depend on $|V_{ub}|$ as illustrated in Figure 1. Such decays are denoted as $B\rightarrow X_{u}\ell\nu$. Despite the experimental challenges due to the neutrino in the final state, [^1] Semitauonic decays are not considered, being even more challenging due to multiple neutrinos in the final state. charmless semileptonic decays offer the best sensitivity to $|V_{ub}|$. Determinations from fully hadronic $B$ meson decays suffer from significant QCD uncertainties [3], while fully leptonic decays of charged $B$ mesons are extremely challenging to study with the available data sets as the light-lepton channels are helicity suppressed and the tau channel suffers from large backgrounds [4].
@@
Figure 1: Feynman diagram of the leading $B\rightarrow X_{u}\ell\nu$ transition. The $X_{u}$ system can either represent a specific meson or a non-resonant system comprising several hadrons produced by the fragmentation of the $u\bar{q}$ system.
@@
The two most recent averages of $|V_{ub}|$ extracted from exclusive and inclusive decays read [5]
$|V_{ub}|_{\rm excl}$ $=(3.43\pm 0.12)\times 10^{-3},$
$|V_{ub}|_{\rm incl}$ $=(4.06\pm 0.16)\times 10^{-3},$ (1)
where the total (experimental and theoretical) uncertainty is given. The two values disagree by about three standard deviations, which limits the precision of the CKM unitarity test. This discrepancy calls for a deeper investigation of the theoretical and experimental factors that may contribute to it. The measurement presented here uses the inclusive method.
The decay width of inclusive charmless semileptonic decays can be described by the heavy quark expansion (HQE) [6, 7], an expansion in powers of the strong coupling constant $\alpha_{s}$ and inverse powers of the $b$ quark mass $m_{b}$. Decay rate predictions in the HQE framework rely on quark-hadron duality which states that predictions made at the parton level, when integrated over sufficiently large portions of the phase space, should yield reliable predictions of hadronic observables [8]. However, selections on kinematic variables such as the lepton energy or the $X_{u}$ mass are necessary in experimental analyses to suppress backgrounds. The most important source of background emerges from CKM-favored semileptonic decays to a hadronic system containing a charm quark. Such decays, denoted as $B\rightarrow X_{c}\ell\nu$, are enhanced by about a factor 50 compared to $B\rightarrow X_{u}\ell\nu$ decays. These kinematic selections break the inclusivity requirement of the HQE [9] and, instead, predictions of differential decay rates are necessary. These predictions in restricted kinematical regions rely on an effective description of the $b$-quark motion inside the meson through a non-perturbative shape function, whose development has been a major theoretical challenge. [^2] Beyond the leading order in the HQE, additional shape functions arise. Treatment of such non-perturbative effects has led to the development of several theoretical frameworks that yield different predictions of the partial decay rate $\Delta\tilde{\Gamma}(B\rightarrow X_{u}\ell\nu)$ ($\Delta\tilde{\Gamma}(B\rightarrow X_{u}\ell\nu)=\Delta\Gamma(B\rightarrow X_{u}\ell\nu)/|V_{ub}|^{2}$) over a given region of phase space [10, 11, 9]. A value of $|V_{ub}|$ can be extracted from the measurement of the inclusive partial branching fraction $\Delta\mathcal{B}(B\rightarrow X_{u}\ell\nu)$ via
@@
In this paper, an inclusive measurement of partial branching fractions of $B\rightarrow X_{u}\ell\nu$ ($\ell=e,\mu$) decays using $e^{+}e^{-}\to\Upsilon(4S)\to B\bar{B}$ data collected by the Belle II experiment is presented. Three variables are used to describe the kinematics of inclusive semileptonic decays: the lepton energy $E_{\ell}$ – with the superscript $B$ added when the energy is calculated in the signal $B$ meson rest frame; the squared momentum transfer defined as $q^{2}=(p_{\ell}+p_{\nu})^{2}$ where $p_{\ell}$ and $p_{\nu}$ are the charged lepton and neutrino four momenta; and the hadronic system mass $M_{X}$. The measurement is performed in three phase-space regions covering about 87%, 57% and 32% of the full $B\rightarrow X_{u}\ell\nu$ phase space and defined by three sets of selections on $E_{\ell}^{B}$, $q^{2}$ and $M_{X}$ designed to suppress the $B\rightarrow X_{c}\ell\nu$ background. For each measured partial branching fraction, $|V_{ub}|$ is extracted from Equation 2 using different decay rate predictions.
@@
The remainder of this manuscript is organized as follows: the collected and simulated data sets used for the measurement are presented in Section II; the event reconstruction and signal selections are outlined in Section III; checks on the modeling of the $B\rightarrow X_{c}\ell\nu$ background and associated corrections are detailed in Section IV; the signal extraction procedure is discussed in Section V; finally, the results of the measurement are presented in Section VI and discussed in Section VII. Throughout this article, charge conjugation is implied and natural units $\hbar=c=1$ are used.
@@ II Data set and simulation
The Belle II detector is situated at the SuperKEKB asymmetric-energy electron–positron collider, based at KEK in Tsukuba, Japan [12]. It is a large solid-angle cylindrical detector consisting of several layers of sub-detectors around the $e^{+}e^{-}$ interaction point (IP). The $z$ axis is defined as the symmetry axis of the detector in the direction of the electron beam, the $y$ axis corresponds to the vertical axis, pointing upwards and the $x$ axis points towards the outside of the accelerator ring. At the innermost part of Belle II is the pixel vertex detector (PXD), which consists of two layers of silicon pixel sensors located at 1.4 cm and 2.2 cm from the IP, which provides precise vertexing capabilities. In the data set used in this measurement, the second pixel layer is installed in only one-sixth of the solid angle. Surrounding the PXD is the silicon vertex detector (SVD), which comprises four layers of silicon strip sensors. The SVD covers polar angles ranging from $17^{\circ}$ to $150^{\circ}$, which further enhances the vertex reconstruction and improves the overall tracking performance. The central drift chamber (CDC) surrounds the SVD and is specifically designed for charged particle tracking. It contains 14,336 sense wires operating in a helium-ethane gas mixture and also covers polar angles ranging from $17^{\circ}$ to $150^{\circ}$. The CDC provides sub-percent accuracy in measuring charged particle momenta, essential for precise kinematic reconstruction. A time-of-propagation (TOP) detector positioned outside the CDC in the barrel region uses Cherenkov light to identify charged particles. In the forward end-cap, this role is fulfilled by a ring-imaging Cherenkov detector equipped with an aerogel radiator (ARICH). Beyond the TOP and ARICH lies the electromagnetic calorimeter (ECL), composed of 8,736 homogeneous thallium-doped caesium iodide crystals. The ECL is divided into three regions: the barrel and two end-caps, and covers polar angles from $12.4^{\circ}$ to $155.1^{\circ}$. This sub-detector excels in detecting electromagnetic showers from both charged and neutral particles and provides percent-level energy resolution for electrons and photons. Additionally, the ECL provides three-momentum information for neutral particles through energy deposition (cluster) analysis. The ECL is surrounded by a superconducting solenoid which generates a magnetic field of 1.5 T. Charged tracks are bent in the $x$-$y$ plane by the magnetic field and the curvature of the trajectory can be used to determine the transverse momentum of the particle. The outermost sub-detector of Belle II is the $K^{0}_{L}$ and muon detector (KLM). The KLM consists of 14 iron layers, each 4.7 cm thick, interspersed with 15 active layers in the barrel region and 14 (12) in the forward (backward) endcap region. The KLM covers polar angles from $20^{\circ}$ to $155^{\circ}$.
@@
The sample of $B\rightarrow X_{u}\ell\nu$ events is a mixture of non-resonant modes (such as $B\to\pi\pi\ell\nu$, $B\to\pi\rho\ell\nu$…), simulated via inclusive $B\rightarrow X_{u}\ell\nu$ frameworks, and resonant modes ($B\rightarrow\pi\ell\nu$, $B\rightarrow\omega\ell\nu$…). The theoretical description of resonant modes is based on functions known as form factors, which encapsulate the non-perturbative dynamics of the transition. Non-resonant and resonant decays are combined using the hybrid model initially proposed in Ref. [18]. After presenting the samples of $B\rightarrow X_{u}\ell\nu$ decays, the hybrid model is described.
The simulation of non-resonant events is based on the De Fazio, Neubert (DFN) model [19] implemented in the EvtGen generator. This model describes the differential $B\rightarrow X_{u}\ell\nu$ rate at order $\mathcal{O}(\alpha_{s})$ and at leading order in the HQE and requires two input parameters for the modeling of events, $m_{b}$ and a non-perturbative parameter $a$ used in the modeling of the shape function. In the Kagan, Neubert renormalization scheme [20] their values read: $m_{b}^{\mathrm{KN}}=4.66\pm 0.04$ GeV and $a^{\mathrm{KN}}=1.3\pm 0.5$ [21]. They are related to the difference between the $B$ meson mass and the $b$ quark mass and the average momentum squared of the $b$ quark inside the $B$ meson. The fragmentation of the hadronic system produced in non-resonant events is simulated by Pythia 8.
In addition, five resonant decays are included in the $B\rightarrow X_{u}\ell\nu$ modeling: $B\to(\pi/\rho/\omega/\eta/\eta^{\prime})\ell\nu$. Their theoretical descriptions rely on specific form factor parametrizations which are also implemented in EvtGen. The Bourrely, Caprini, Lellouch [22] parametrization is used to model $B\rightarrow\pi\ell\nu$ decays. The form factor parameters are taken from Table 57 in Ref. [23]; $B\rightarrow\rho/\omega\ell\nu$ decays are modelled with the Bharucha, Straub, Zwicky parametrization [24] whose values are taken from Table 4 in Ref. [25]; $B\rightarrow\eta/\eta^{\prime}\ell\nu$ decays are modelled via the Isgur, Scora, Grinstein, Wise approach [26] with input parameters taken from the light-cone sum rules framework described in Ref. [27].
Resonant and non-resonant $B\rightarrow X_{u}\ell\nu$ contributions are combined using the hybrid approach. The resonant and non-resonant contributions are summed in three-dimensional bins of $E_{\ell}^{B}$, $M_{X}$ and $q^{2}$. In each bin, the inclusive contribution ($\Delta\mathcal{B}^{\textrm{inc}}$) is obtained from the non-resonant component and the exclusive contribution ($\Delta\mathcal{B}^{\textrm{exc}}$) is obtained from the sum of exclusive components. The non-resonant part is scaled down such that the sum of non-resonant and resonant contributions matches the inclusive partial branching fraction. The non-resonant contribution is therefore scaled by weights defined as:
@@
where the indices $i,j,k$ denote a three-dimensional bin. These three-dimensional bins are defined by the combinations of the following one-dimensional bins:
$E_{\ell}^{B}$ $:[0,0.50,1.00,1.25,1.50,1.75,2.00,2.25,3.00]\rm~GeV,{\rm~and}$
$q^{2}$ $:[0,2.50,5.00,7.50,10.0,12.5,15.0,20.0,25.0]\rm~GeV^{2},$
$M_{X}$ $:[0,1.40,1.60,1.80,2.00,2.50,3.00,3.50]\rm~GeV.$
The $B\rightarrow X_{c}\ell\nu$ background is modelled as a sum of different types of resonant decays. Each resonance is modelled with a specific form factor parametrization implemented in EvtGen. The $B\rightarrow D\ell\nu$ and $B\rightarrow D^{*}\ell\nu$ decays are modelled using the Bernlochner, Ligeti, Papucci, Robinson, Xiong, Prim form factor model [28] with eight common parameters. The four heavy $1P$ states $D_{0}^{*}$, $D_{1}^{\prime}$, $D_{1}$ and $D_{2}^{*}$ are collectively referred to as $D^{**}$ [29]. The $B\rightarrow D^{**}\ell\nu$ decays are modelled using the Bernlochner, Ligeti, Robinson (BLR) model from Ref. [30] with three parameters for the broad $D_{0}^{*}$ and $D_{1}^{\prime}$ states and four parameters for the narrower $D_{1}$ and $D_{2}^{*}$ states. The $B\rightarrow X_{c}\ell\nu$ modeling also includes the decays $B\to D^{(*)}\pi\pi\ell\nu$ and $B\to D_{s}^{(*)}K\ell\nu$. The $B\to D^{(*)}\eta\ell\nu$ modes are simulated to fill the gap between the inclusively measured $B\rightarrow X_{c}\ell\nu$ rate and the sum of exclusive decay rates [31]. The $D^{(*)}\pi\pi$ and $D^{(*)}\eta$ modes are collectively referred to as gap modes and are denoted $B\rightarrow D^{**}_{\textrm{Gap}}\ell\nu$. As there exists no theoretical description of these decays, they are simulated via intermediate decays to $D^{**}$ resonances using the BLR form factor parametrization. The $D\eta$ decays are simulated via $B\to D_{0}^{*}(\to D\eta)\ell\nu$ and $D^{*}\eta$ decays via $B\to D_{1}^{\prime}(\to D^{*}\eta)\ell\nu$. The $D^{(*)}\pi\pi$ modes are equally split between intermediate decays to $D_{0}^{*}$ and $D_{1}^{\prime}$.

### v2

# Belle II Collaboration

@@ Abstract
A sample of 365 fb-1 of $e^{+}e^{-}\to\Upsilon(4S)\to B\bar{B}$ data collected by the Belle II experiment is used to measure the partial branching fractions of charmless semileptonic $B$ meson decays and determine the magnitude of the Cabibbo-Kobayashi-Maskawa matrix element $V_{ub}$. Events containing a signal electron or muon $\ell$ and a fully reconstructed hadronic $B$ decay that constrains the signal kinematics are selected, while the rest of the event defines the hadronic system $X_{u}$ associated with the signal. To discriminate the signal from the 50-times larger background originating from CKM-favored semileptonic $B$ decays, a template fit is performed in both signal and control regions after applying an optimized selection. The partial branching fraction measured for lepton energies greater than 1 GeV in the signal $B$ meson rest frame is $\Delta\mathcal{B}(B\rightarrow X_{u}\ell\nu)=(1.54\pm 0.08\,{\rm(stat.)}\pm 0.12\,{\rm(syst.)})\times 10^{-3}$. From this measurement, using the Gambino, Giordano, Ossola, Uraltsev theoretical framework, $|V_{ub}|=(4.01\pm 0.19^{+0.07}_{-0.08})\times 10^{-3}$ is determined, where the uncertainties are experimental and theoretical, respectively. This value is consistent with the world average obtained from previous inclusive measurements. Different theoretical predictions and partial branching fractions measured in other phase-space regions, defined by additional selections on the $X_{u}$ and leptonic system masses, are also used to determine $|V_{ub}|$. This allows for a comparison of the resulting values across theoretical frameworks and phase-space regions.
@@ I Introduction
Precise measurements of the Cabibbo-Kobayashi-Maskawa (CKM) matrix elements [26, 40] are crucial to test the quark weak interactions described by the standard model of particle physics. For three fermion generations the CKM matrix is unitary, which imposes strict relations between its elements. The precision of experimental tests of several of these relations is limited by the current uncertainty in the determination of the strength of the $b$-to-$u$ quark coupling, $|V_{ub}|$. $B$-meson decays to either an electron or a muon, $\ell$, a neutrino, $\nu$, and a charmless hadronic system, $X_{u}$, depend on $|V_{ub}|$ as illustrated in Fig. 1. Such decays are denoted as $B\rightarrow X_{u}\ell\nu$. Despite the experimental challenges due to the neutrino in the final state, [^1] Semitauonic decays are not considered, being even more challenging due to multiple neutrinos in the final state. charmless semileptonic decays offer the best sensitivity to $|V_{ub}|$. Determinations from fully hadronic $B$ meson decays suffer from significant QCD uncertainties [1], while fully leptonic decays of charged $B$ mesons are extremely challenging to study with the available datasets as the light-lepton channels are helicity suppressed and the tau channel suffers from large backgrounds [4].
@@
Figure 1: Feynman diagram of the leading $B\rightarrow X_{u}\ell\nu$ transition. The $X_{u}$ system can either represent a specific meson or a nonresonant system comprising several hadrons produced by the fragmentation of the $u\bar{q}$ system.
@@
[…]
$V_{ub}$—_incl = (4.06 ±0.16) ×10^-3, where the total (experimental and theoretical) uncertainty is given. The two values disagree by about three standard deviations, which limits the precision of the CKM unitarity test. This discrepancy calls for a deeper investigation of the theoretical and experimental factors that may contribute to it. The measurement presented here uses the inclusive method.
The decay width of inclusive charmless semileptonic decays can be described by the heavy quark expansion (HQE) [48, 28], an expansion in powers of the strong coupling constant $\alpha_{s}$ and inverse powers of the $b$ quark mass $m_{b}$. Decay rate predictions in the HQE framework rely on quark-hadron duality which states that predictions made at the parton level, when integrated over sufficiently large portions of the phase space, should yield reliable predictions of hadronic observables [22]. However, selections on kinematic variables such as the lepton energy or the $X_{u}$ mass are necessary in experimental analyses to suppress backgrounds. The most important source of background emerges from CKM-favored semileptonic decays to a hadronic system containing a charm quark. Such decays, denoted as $B\rightarrow X_{c}\ell\nu$, are enhanced by about a factor 50 compared to $B\rightarrow X_{u}\ell\nu$ decays. These kinematic selections break the inclusivity requirement of the HQE [35] and, instead, predictions of differential decay rates are necessary. These predictions in restricted kinematical regions rely on an effective description of the $b$-quark motion inside the meson through a nonperturbative shape function, whose development has been a major theoretical challenge. [^2] Beyond the leading order in the HQE, additional shape functions arise. Treatment of such nonperturbative effects has led to the development of several theoretical frameworks that yield different predictions of the partial decay rate $\Delta\tilde{\Gamma}(B\rightarrow X_{u}\ell\nu)$ ($\Delta\tilde{\Gamma}(B\rightarrow X_{u}\ell\nu)=\Delta\Gamma(B\rightarrow X_{u}\ell\nu)/|V_{ub}|^{2}$) over a given region of phase space [42, 9, 35]. A value of $|V_{ub}|$ can be extracted from the measurement of the inclusive partial branching fraction $\Delta\mathcal{B}(B\rightarrow X_{u}\ell\nu)$ via
@@
In this paper, an inclusive measurement of partial branching fractions of $B\rightarrow X_{u}\ell\nu$ ($\ell=e,\mu$) decays using $e^{+}e^{-}\to\Upsilon(4S)\to B\bar{B}$ data collected by the Belle II experiment is presented. Three variables are used to describe the kinematics of inclusive semileptonic decays: the lepton energy $E_{\ell}$ – with the superscript $B$ added when the energy is calculated in the signal $B$ meson rest frame; the squared momentum transfer defined as $q^{2}=(p_{\ell}+p_{\nu})^{2}$ where $p_{\ell}$ and $p_{\nu}$ are the charged lepton and neutrino four momenta; and the hadronic system mass $M_{X}$. The measurement is performed in three phase-space regions covering about 87%, 57% and 32% of the full $B\rightarrow X_{u}\ell\nu$ phase space and defined by three sets of selections on $E_{\ell}^{B}$, $q^{2}$ and $M_{X}$ designed to suppress the $B\rightarrow X_{c}\ell\nu$ background. For each measured partial branching fraction, $|V_{ub}|$ is extracted from Eq. 1 using different decay rate predictions.
@@
The remainder of this manuscript is organized as follows: the collected and simulated datasets used for the measurement are presented in Sec. II; the event reconstruction and signal selections are outlined in Sec. III; checks on the modeling of the $B\rightarrow X_{c}\ell\nu$ background and associated corrections are detailed in Sec. IV; the signal extraction procedure is discussed in Sec. V; finally, the results of the measurement are presented in Sec. VI and discussed in Sec. VII. Throughout this article, charge conjugation is implied and natural units $\hbar=c=1$ are used.
@@ II Dataset and simulation
The Belle II detector is situated at the SuperKEKB asymmetric-energy electron–positron collider, based at KEK in Tsukuba, Japan [6]. It is a large solid-angle cylindrical detector consisting of several layers of subdetectors around the $e^{+}e^{-}$ interaction point (IP). The $z$ axis is defined as the symmetry axis of the detector in the direction of the electron beam, the $y$ axis corresponds to the vertical axis, pointing upward and the $x$ axis points towards the outside of the accelerator ring. At the innermost part of Belle II is the pixel vertex detector (PXD), which consists of two layers of silicon pixel sensors located at 1.4 cm and 2.2 cm from the IP, which provides precise vertexing capabilities. In the dataset used in this measurement, the second pixel layer is installed in only one-sixth of the solid angle. Surrounding the PXD is the silicon vertex detector (SVD), which comprises four layers of silicon strip sensors. The SVD covers polar angles ranging from $17^{\circ}$ to $150^{\circ}$, which further enhances the vertex reconstruction and improves the overall tracking performance. The central drift chamber (CDC) surrounds the SVD and is specifically designed for charged particle tracking. It contains 14,336 sense wires operating in a helium-ethane gas mixture and also covers polar angles ranging from $17^{\circ}$ to $150^{\circ}$. The CDC provides subpercent accuracy in measuring charged particle momenta, essential for precise kinematic reconstruction. A time-of-propagation (TOP) detector positioned outside the CDC in the barrel region uses Cherenkov light to identify charged particles. In the forward end-cap, this role is fulfilled by a ring-imaging Cherenkov detector equipped with an aerogel radiator (ARICH). Beyond the TOP and ARICH lies the electromagnetic calorimeter (ECL), composed of 8,736 homogeneous thallium-doped caesium iodide crystals. The ECL is divided into three regions: the barrel and two end-caps, and covers polar angles from $12.4^{\circ}$ to $155.1^{\circ}$. This subdetector excels in detecting electromagnetic showers from both charged and neutral particles and provides percent-level energy resolution for electrons and photons. Additionally, the ECL provides three-momentum information for neutral particles through energy deposition (cluster) analysis. The ECL is surrounded by a superconducting solenoid which generates a magnetic field of 1.5 T. Charged tracks are bent in the $x$-$y$ plane by the magnetic field and the curvature of the trajectory can be used to determine the transverse momentum of the particle. The outermost subdetector of Belle II is the $K^{0}_{L}$ and muon detector (KLM). The KLM consists of 14 iron layers, each 4.7 cm thick, interspersed with 15 active layers in the barrel region and 14 (12) in the forward (backward) endcap region. The KLM covers polar angles from $20^{\circ}$ to $155^{\circ}$.
@@
The sample of $B\rightarrow X_{u}\ell\nu$ events is a mixture of nonresonant modes (such as $B\to\pi\pi\ell\nu$, $B\to\pi\rho\ell\nu$…), simulated via inclusive $B\rightarrow X_{u}\ell\nu$ frameworks, and resonant modes ($B\rightarrow\pi\ell\nu$, $B\rightarrow\omega\ell\nu$…). The theoretical description of resonant modes is based on functions known as form factors, which encapsulate the nonperturbative dynamics of the transition. Nonresonant and resonant decays are combined using the hybrid model initially proposed in Ref. [56]. After presenting the samples of $B\rightarrow X_{u}\ell\nu$ decays, the hybrid model is described.
The simulation of nonresonant events is based on the De Fazio, Neubert (DFN) model [30] implemented in the EVTGEN generator. This model describes the differential $B\rightarrow X_{u}\ell\nu$ rate at order $\mathcal{O}(\alpha_{s})$ and at leading order in the HQE and requires two input parameters for the modeling of events, $m_{b}$ and a nonperturbative parameter $a$ used in the modeling of the shape function. In the Kagan, Neubert renormalization scheme [38] their values read: $m_{b}^{\mathrm{KN}}=4.66\pm 0.04$ GeV and $a^{\mathrm{KN}}=1.3\pm 0.5$ [25]. They are related to the difference between the $B$ meson mass and the $b$ quark mass and the average momentum squared of the $b$ quark inside the $B$ meson. The fragmentation of the hadronic system produced in nonresonant events is simulated by PYTHIA 8.
In addition, five resonant decays are included in the $B\rightarrow X_{u}\ell\nu$ modeling: $B\to(\pi/\rho/\omega/\eta/\eta^{\prime})\ell\nu$. Their theoretical descriptions rely on specific form factor parametrizations which are also implemented in EVTGEN. The Bourrely, Caprini, Lellouch [24] parametrization is used to model $B\rightarrow\pi\ell\nu$ decays. The form factor parameters are taken from Table 57 in Ref. [10]; $B\rightarrow\rho/\omega\ell\nu$ decays are modeled with the Bharucha, Straub, Zwicky parametrization [20] whose values are taken from Table 4 in Ref. [19]; $B\rightarrow\eta/\eta^{\prime}\ell\nu$ decays are modeled via the Isgur, Scora, Grinstein, Wise approach [57] with input parameters taken from the light-cone sum rules framework described in Ref. [31].
Resonant and nonresonant $B\rightarrow X_{u}\ell\nu$ contributions are combined using the hybrid approach. The resonant and nonresonant contributions are summed in three-dimensional bins of $E_{\ell}^{B}$, $M_{X}$ and $q^{2}$. In each bin, the inclusive contribution ($\Delta\mathcal{B}^{\textrm{inc}}$) is obtained from the nonresonant component and the exclusive contribution ($\Delta\mathcal{B}^{\textrm{exc}}$) is obtained from the sum of exclusive components. The nonresonant part is scaled down such that the sum of nonresonant and resonant contributions matches the inclusive partial branching fraction. The nonresonant contribution is therefore scaled by weights defined as:
@@
where the indices $i,j,k$ denote a three-dimensional bin. These three-dimensional bins are defined by the combinations of the following one-dimensional bins: {align*} E_ℓ^B : [0, 0.50, 1.00, 1.25, 1.50, 1.75, 2.00, 2.25, 3.00]  GeV,  and
q^2 : [0, 2.50, 5.00, 7.50, 10.0, 12.5, 15.0, 20.0, 25.0]  GeV ^2,
M_X : [0, 1.40, 1.60, 1.80, 2.00, 2.50, 3.00, 3.50]  GeV.
The $B\rightarrow X_{c}\ell\nu$ background is modeled as a sum of different types of resonant decays. Each resonance is modeled with a specific form factor parametrization implemented in EVTGEN. The $B\rightarrow D\ell\nu$ and $B\rightarrow D^{*}\ell\nu$ decays are modeled using the Bernlochner, Ligeti, Papucci, Robinson, Xiong, Prim form factor model [16] with eight common parameters. The four heavy $1P$ states $D_{0}^{*}$, $D_{1}^{\prime}$, $D_{1}$ and $D_{2}^{*}$ are collectively referred to as $D^{**}$ [18]. The $B\rightarrow D^{**}\ell\nu$ decays are modeled using the Bernlochner, Ligeti, Robinson (BLR) model from Ref. [17] with three parameters for the broad $D_{0}^{*}$ and $D_{1}^{\prime}$ states and four parameters for the narrower $D_{1}$ and $D_{2}^{*}$ states. The $B\rightarrow X_{c}\ell\nu$ modeling also includes the decays $B\to D^{(*)}\pi\pi\ell\nu$ and $B\to D_{s}^{(*)}K\ell\nu$. The $B\to D^{(*)}\eta\ell\nu$ modes are simulated to fill the gap between the inclusively measured $B\rightarrow X_{c}\ell\nu$ rate and the sum of exclusive decay rates [27]. The $D^{(*)}\pi\pi$ and $D^{(*)}\eta$ modes are collectively referred to as gap modes and are denoted $B\rightarrow D^{**}_{\textrm{Gap}}\ell\nu$. As there exists no theoretical description of these decays, they are simulated via intermediate decays to $D^{**}$ resonances using the BLR form factor parametrization. The $D\eta$ decays are simulated via $B\to D_{0}^{*}(\to D\eta)\ell\nu$ and $D^{*}\eta$ decays via $B\to D_{1}^{\prime}(\to D^{*}\eta)\ell\nu$. The $D^{(*)}\pi\pi$ modes are equally split between intermediate decays to $D_{0}^{*}$ and $D_{1}^{\prime}$.